OpenStudy (narissa):

please help solve using substitution method d+e=2 d-e=4

OpenStudy (narissa):

@HossamHoussien

OpenStudy (neonumbrella5115):

Do you know what the substitution method is? Do you have a guess?

OpenStudy (narissa):

no solution,infinite number of solutions,solution is 5,1 or solution is 3,-1

OpenStudy (neonumbrella5115):

Do you know what the substitution method is?

OpenStudy (narissa):

no solution?

OpenStudy (narissa):

yes i do

OpenStudy (narissa):

i need help thats y im here

OpenStudy (neonumbrella5115):

Okay, no solution is not the answer. To do substitution, first you need to solve each equation for e. Can you do that?

OpenStudy (narissa):

ok 3,-1

OpenStudy (neonumbrella5115):

Yes that is the correct answer! Good job!

OpenStudy (hossamhoussien):

Look here it asking to use substitution method which is express one variable in terms of another variable Ex: \[y=x+2\] \[y+2=-4\] So what you have to do is to remove each y and plug (x+2) instead \[(x+2)+2=-4\] \[x+4=-4\] \[x=-8\] and then plug x-value into very first equation to get y-value

OpenStudy (hossamhoussien):

try to follow these steps in your question and if you need any help i'm here